Part A

50 grams of ice-cold water (0 °C) is heated to boiling (100 °C). The specific heat capacity of water is 4.18 J /g°C. How much energy did the water absorb?


Part B

Over the next hour, the water gradually cools to 25 °C. How much energy did the water lose?


Part C

What was the temperature of the room that the water was in?


1
Expert's answer
2021-11-16T15:44:03-0500

Part A

"Q = mcΔT \\

m= 50\; g \\

ΔT = 100 -0 = 100 \;°C \\

Q = 50 \times 4.18 \times 100 = 20900 \;J"

Part B

"ΔT = 25-100 = -75 \; °C \\

Q = 50 \times 4.18 \times (-75) = -15675 \; J"

the water lose 15675 J

Part C

The temperature of the room was 25 °C